OpenStudy (anonymous):

Which point is NOT a solution of the inequality x + 2y ≤ 8? Answer key: (0, 4) (4, 4) (-2, 3) (6, 0)

OpenStudy (mertsj):

Why don't you substitute in the given points until you find one that makes a false statement. I will do the first one for you: \[0+2(4)\le 8\] \[8\le 8\] That is true since 8 is equal to 8

OpenStudy (mertsj):

Now you try the others.

OpenStudy (anonymous):

so Im not sure if its correct but I think its (6,0)???????

OpenStudy (mertsj):

\[6+2(0)\le 8\] \[6\le 8\] That says that 6 is less than or equal to 8. Since 6 IS less than 8, that is true.

OpenStudy (anonymous):

thankyou

OpenStudy (mertsj):

Did you try the others? Remember, you are looking for the FALSE one.

OpenStudy (anonymous):

-2,3 ?

OpenStudy (anonymous):

??